a student dissolves 0.461g of potassium hydrogen phthalate (KHP) solution with a solution hydroxide. She records her initial burst reading as 4.72 mL and her final burst reading as 14.26mL. what volume sodium hydroxide did she use? how many moles of KHP were present in the solution? if one mole of khp reacts with one mole of hydroxide, how many moles of hydroxide were added?

what was the molarity of the sodium hydroxide solution?
Answer #1

In the given reaction, one mole of potassium hydrogen phthalate (KHP) reacts with the one mole of NaOH.

Initial burst reading of burette = 4.72 ml

final burst reading of burette = 14.26 ml

The volume used of NaOH = 14.26 - 4.72 = 9.54 mL

Molar mass of KHP = 204.22 g/mol

Number of moles of KHP in 0.461 g = 0.461/204.22 = 0.00225 mol

Since, final burst reading observed when number of moles of KHP reacts with same number of moles of NaOH,

I.e., number of moles of NaOH = 0.00225 mole

Molarity of hydroxide = number of moles/liter

or,